### v1

# First-order and Berezinskii-Kosterlitz-Thouless phase transitions in two-dimensional generalized XY models

Abstract
The aim of this paper is to illustrate that generalized two-dimensional XY models (proposed by Romano and Zagrebnov) may also support a first-order phase transition. Two approaches are employed to accurately determine the critical parameter $q$ at which such a transition takes place. Furthermore, we show that the model is characterized by three distinct regions concerning both first-order and Berezinskii-Kosterlitz-Thouless phase transitions. Finally, the underlying mechanisms governing such transitions are presented, along with an estimation of the critical temperatures.
preprint: APS/123-QED
I introduction
Figure 1: (a) Energy for spin in function of temperature for different values of q, and $L=96$. We can observe a discontinuity being formed with the increase of q.Specific heat for q=3 (b), and the scaling of max specific heat with $L^{2}$ (c), showing comportament of a first order transition.
Two-dimensional magnetic models characterized by short-range interaction and continuous symmetry fail to exhibit long-range order at a finite temperature [1]. Despite the absence of long-range order due to the destructive influence of spin waves, even at low temperatures, a quasi-long-range order with a power-law decaying correlation function is evident. At higher temperatures, correlations decay exponentially, leading to a disordered phase. The transition from quasi-long-range order to a disordered phase constitutes a topological phase transition within the Berezinskii–Kosterlitz–Thouless (BKT) universality class [2, 3]. This transition is characterized by the unbinding of topological objects, know as vortices, at a critical temperature $T_{BKT}$.
Among the simplest models supporting topological excitations in the spin field and a BKT transition are those described by the Hamiltonian $H=-J\sum_{<ij>}(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j})$, where $J$ is a coupling constant, the summation is over nearest-neighbors sites in a square lattice, and $S^{\alpha}_{i}$ represents the spin at site $i$. From this Hamiltonian one can extract the Planar Rotator Model (PRM) by enforcing the constraint $S^{2}_{x}+S^{2}_{y}=1$, which confines the spins to the plane. On the other hand, this same Hamiltonian leads to the XY-model, simply by allowing spins to have three components and the constraint $S^{2}_{x}+S^{2}_{y}+S^{2}_{z}=1$. Both models can be parameterized by using scalar fields. For the PRM, the azimuthal angle $\phi$ is enough, yielding the Hamiltonian $H=-j\sum_{<ij>}\cos(\phi_{i}-\phi_{j})$. Meanwhile, for the XY model, both azimuthal and polar angles are employed, resulting in the Hamiltonian $H=-j\sum_{<ij>}\sin\theta_{i}\sin\theta_{j}\cos(\phi_{i}-\phi_{j})$.
Extensive studies on the static critical behavior of such models have established consensus regarding phase transitions, critical temperature and exponents [4] [5]. The nature of the phase transitions was very investigated in the 1970’s and 1980’s, such that different generalizations and extensions of the these two models have become also object of studies.
Domany et al. [6] introduced a generalization of the PRM model utilizing the Villain form $V(\phi)=1-\cos^{p^{2}}(\phi^{\prime}/2)$, with $\phi^{\prime}=\phi_{i}-\phi_{j}$. As $p$ increases, a narrow potential well with width $\pi/p$ emerges, resembling the $n$-states Potts model with $p$ proportional to $n$. With Monte Carlo simulations, Domany et al. [6] revealed that an increase in $p$ leads to a discontinuity in energy and a two-peak structure in the energy histogram, suggesting a first-order transition for large $p$. Some studies challenged this result, particularly those based on renormalization group analysis [7, 8]. However, the possibility of first-order induced disorder in the planar model was demonstrated [9], and later, rigorous proof by van Enter and Shlosman confirmed that $SO(n)$ invariant $n$-vector models with sufficiently deep and narrow minima can undergo a first-order phase transition[10].
A generalized XY-model was proposed by Romano and Zagrebnov[11], whose Hamiltonian is given by,
$H^{Gen}_{XY}=$ $-J\sum_{<i,j>}\big{[}1-(S^{z}_{i}S^{z}_{i})^{2}-(S^{z}_{j}S^{z}_{j})^{2}$ (1)
$+(S^{z}_{i}S^{z}_{j})^{2}\big{]}^{(q-1)/2}\left(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}\right).$
Writing this with the help of the two scalar fields $\theta$ and $\phi$, one gets,
$H^{Gen}_{XY}=-J\sum_{<i,j>}(\sin\theta_{i}sin\theta_{j})^{q}\cos\phi_{i}\cos\phi_{j}.$ (2)
Here, $q\in$ N is the generalization parameter, and for $q=1$, the XY model is recovered. In their investigations[11], Romano and Zagrebnov used rigorous inequalities for all values of $q$, considering systems in two and three dimensions. For the two-dimensional (2D) case, they showed that, for arbitrary $q$, the model has orientational disorder at all finite temperatures, undergoing a transition to a low temperature phase with slow decay of correlations and infinite susceptibility. So the class of universality of the transition is not necessarily the same as that of the proper BKT transition. Mól et al. [12] extensively studied the vortex-like solutions within the generalized XY model, employing both the continuous Hamiltonian and the Self Consistent Harmonic Approximation (SCHA) [13]. They showed that only static planar-vortex configurations are stable and supported for any $q\geq 1$, and that the critical temperature of phase transitions decreases with the parameter $q$. In addition, Monte Carlo simulations [14] suggest a first-order phase transition in the generalized XY model for lower values of $q$, but the result was not conclusive. Although the first order transition was initially contested, a rigorous proof by van Enter et al. [15] confirmed that the 2D generalized XY model proposed by Romano and Zagrabenov indeed exhibits a first-order phase transition, leaving open the value of the critical parameter $q_{c}$ at which this transition begins to appear.
Our motivations for the present work are threefold: firstly, we would like to confirm the existence of a first-order phase transition for Hamiltonian (1); the second motivation is to obtain the generalized critical parameter $q_{c}$ where the system exhibits a first-order transition; finally, our task concentrates in providing a physical interpretation for this first-order phase transition. For these purposes, we have used the hybrid Monte Carlo algorithm, not much different from that used by Mól et al. [14], which will be explained briefly in the next section. To enhance readability, most thermodynamic results and techniques for determining transition temperatures are presented in appendices and only the main results are shown in the usual text. We begin by presenting Monte Carlo simulation results and then, guided by these calculations and the solution of van Enter et al. [15], we analytically estimate the the critical temperature of the first-order phase transition.
II Monte Carlo Method
Figure 2: (a) show how the gap varies with the size of the system, comparing the two sizes of L=80,96.(b) show how the transition varies with the temperature for q=3.(c) Lee-Kosterlitz criterion for q=3,5,6, where the expected comportament of a first order transition is observed for the three cases.
We have employed a Hybrid Monte Carlo approach to generate configurations for calculating thermal averages. A Monte Carlo Step (MCS) involves a sequence of operations, starting with a Wolff cluster algorithm for in-plane components [16], followed by modified Metropolis single-spin updates for the three components and $N$ over-relaxation steps. In the modified Metropolis algorithm [17], a randomly selected spin undergoes a small increment in random components, followed by renormalization to maintain unit length. The acceptance is determined by the standard Metropolis algorithm, with the spin increment length chosen to achieve an acceptance rate between 30% and 60%. This modification proved effective in mitigating critical slowdown at low temperatures. Various acceptance rates were tested, with little impact on the results. Over-relaxation involved reflecting a randomly selected spin across the effective magnetic field due to its neighbors while preserving the spin length. This algorithm effectively alters configurations while maintaining energy conservation. Mean values were obtained using 10,000 Monte Carlo passes to warm up the samples, and 300,000 configurations were individually divided into each bin for analysis. Testing higher numbers of configurations and bin sizes showed minimal impact on the results.
III Results
III.1 Monte Carlo Results
Initially, we will examine the impact of the generalized parameter $q$ on the energy per spin ($e$) and the specific heat ($c_{v}$). For the simulations involving the usual XY-model ($q=1$), our results align with expectations [18]. Both energy and specific heat exhibit minor finite-size effects as we approach the critical temperature; the estimated errors escalate due to rising fluctuations and critical slowing down. A distinct peak emerges at a temperature exceeding the critical temperature of the BKT transition. The determination of this critical temperature involves various methods, detailed in table 1 and further explicated in the appendix. With the increasing of the parameter $q$, the specific heat peak moves toward lower temperatures (still above the critical temperature observed for a BKT transition) and becomes narrow and higher as the system size $L$ increases. The energy begins to exhibit a discontinuity (Fig 1a) at the same temperature where the maximum specific heat ($c_{v}^{max}$) occurs. These two characteristics hint at the possibility of a first-order phase transition. However, to definitively characterize a first-order transition from simulations, it is imperative to observe that the transition persists in the thermodynamic limit ($L\rightarrow\infty$). Regarding the specific heat [19], the maximum value exhibits a proportional relationship with the volume of the system dimension, expressed as $c_{v}^{max}\propto L^{d}$. Therefore, plotting $c_{v}^{max}$ versus $L^{2}$ for various values of $q$ can reveal if they present a linear dependence. As Fig.1c shows, for $q=1$, no finite-size effect is observed. However, for $q=3$ and higher, the specific heat peak continues to increase with the size of the system, but obtaining the precise maximum value becomes increasingly challenging due to fluctuations.
A more effective approach for discerning a first-order phase transition involves computing the histogram of the energy distribution [20]. This is achieved through extended runs within the temperature region surrounding the energy discontinuity for different lattice sizes. In a first-order transition, the theory predicts a double-peak structure [21], each peak representing a distinct state of the system: one ordered, the other disordered. The transition temperature is defined when two peaks with equal areas emerge in the histogram, indicating equiprobability between the two states. For any $q<3$, no double-peak structure was identified for any lattice size $L$, leading to the conclusion that the only one phase transition for the system is of the BKT type. However, for $q=3$, a subtle double-peak histogram is observed, particularly for $L=64$, with a temperature transition difference of less than 1% compared to the energy discontinuity and the specific heat peak temperature. As $L$ increases, the double-peak structure becomes more apparent as shown in 2, because the free energy barrier between the two states becomes higher. The double-peak structure becomes more pronounced with increasing $L$ (see Fig. 2), reflecting a growing free energy barrier between the two states and, again, the transition temperature differs by less than 1% from the energy discontinuity and specific heat peak temperatures.
To definitively confirm the first-order transition, we employ the Lee-Kosterlitz criterion [22]. According to this reasoning, determining a first-order transition involves calculating the free energy barrier, denoted as $\triangle F(L)$, for various sizes of the system $L$. The Lee-Kosterlitz criterion implies that $\triangle F(L)$ should be independent of $L$, while for a first-order phase transition, it is expected to be an increasing function of $L$. So, it is crucial to highlight the dependence on the size of the lattice. For $q=3$, the double-peak structure shows up faintly for $L=64$. For higher values of $q$, the double-peak structure manifests much earlier; for instance, in $q=4$ one can observe it for $L=32$. This inherent size dependence poses challenges in applying the Lee-Kosterlitz criterion to small values of $q$ and the same problem was observed by Lee-Kosterlitz [22] in their own paper for small values of potts model. They attributed this issue to the weakness of the transition. However, for $q=3,5,6$, we observe the anticipated behavior indicative of a first-order phase transition, as depicted in Fig. 2c. Despite computational limitations prevent the acquisition of the Lee-Kosterlitz criterion, the observed increasing in the free energy barrier supports the assertion that the critical parameter is $q=3$. In Table 1, we show the values of the transition temperature utilizing the Binder Fourth Cumulant (see Complementary material). This method reveals another characteristic of a first-order phase transition: a jump to a negative value precedes the attainment of the value $U_{L}=0.5$ for an ordered system [23].
In the range from $q=3$ to $q=5$, the critical temperature obtained through finite-size scaling of the susceptibility and the Helicity ( see complementary material) consistently reveals transition temperatures higher than the actual transition temperature. However, for $q=6$ and higher, the critical temperature and the transition temperature obtained from the Binder Fourth Cumulant and the susceptibility converge. Consequently, the determination of the critical temperature becomes challenging based on the finite-size scaling of the susceptibility and Binder Fourth Cumulant. This observation potentially suggests the absence of the BKT transition for this generalized parameter.
The only one method that has shown possible to determine a critical temperature is through the finite-size scaling (FSS) of the helicity modulus. This is due to the system not only exhibiting vortices but also becoming progressively populated with the increase of the parameter $q$, as illustrated in 3.
Figure 3: Vortex Density in function of temperature for different values of q, and $L=96$. We can observe a discontinuity being formed with the increase of q.
The vortex density exhibits a discontinuity with increasing $q$ at the same temperature as the energy and where the specific heat presents a peak, indicating a correlation between these events. To illustrate this correlation more effectively, we calculate the vortex density difference. Essentially, at the system transitions between states, the vortex density difference jumps from 00 to a non-zero value. To better demonstrate this point, we have taken the vortex density difference as an order parameter, that is, when the system change from one state for another, the vortex density jumps from 00 to a specified value. This fact allowed us to define a way to obtain the critical generalization parameter of the model, simply by calculating the mean vortex density difference for different $q$ at the region of temperature where supposedly the discontinuity should appear. The first value in which the mean vortex density difference is appreciable is found to be $q=3$, agreeing with the result obtained by the histogram and Binder Cumulant.
(a) ”Ordered Phase”
(b) q=6 Histogram
(c) ”Disordered Phase”
Figure 4: Two distinct configurations were observed for $q=6$ at the system’s transition temperature, $T_{c}=0.6212$, with a lattice size of $L=96$. The red points denote vortices, while black points represent antivortices. The state labeled as ”Disordered” with higher energy, exhibits a significantly larger population of vortices compared to the ”Ordered” state with lower energy.
III.2 Gibbs Measures and First Order Transitions
The Gibbs Measures method is a powerful tool for studying first-order phase transitions. For every Hamiltonian $H^{\Lambda}_{\epsilon}(\phi,\theta)$, there exists a Gibbs measure $\mu^{\Lambda}(d\phi,d\theta)$. If this measure is not unique, the system exhibits a first-order transition. A solid background on Gibbs measures and phase transitions can be found in [24, 25]. One of the early contributions using Gibbs Measures was made by Dobrushin and Shlosman [26], where they established a theorem stating that every state of Gibbs on a two-dimensional lattice with a continuous, inferiorly bounded, translationally invariant and rapidly decreasing potential is always an invariant measure under the action of a symmetry group G.
While these conditions may appear restrictive, they are not impossible to satisfy. Indeed, the Heisenberg model meets these conditions, and this theorem essentially extends the Mermin-Wagner theorem, wherein invariance to the potential means the absence of continuous symmetry breaking in two-dimensional systems. The possibility of phase transitions in two-dimensional models with continuous symmetries was further explored by Shlosman in another paper[27], utilizing two powerful methods: The chessboard estimate and the reflection positivity [28, 29, 30, 31]. Shlosman demonstrated that obtaining two phases for isotropic short-range interaction models with continuous symmetry is possible, provided the temperature is low enough. However, all phases remain invariant under an action called G in the configuration space, meaning there is a likelihood of a first-order phase transition without breaking symmetry. As a practical example, Van Enter and Shlosman[10] proved this for the Domany model[6].
Our focus here lies on the results obtained by Van Enter et al.[15] for the XY generalized model. They employed a square-ditch approximation, where only spins within the same ditch can interact. The present implies that a small interval $[\pi/2-\epsilon,\pi/2+\epsilon]$, is defined for the polar angle $\theta$, where the ditch parameter $\epsilon$ is a function of the $q$-parameter, i.e. $\epsilon(q)$. In addition, $\epsilon(q)$ must decay at least with $1/\sqrt{q}$. This approximation implies that two spins interact only when they are in the same ditch. The aforementioned allows them, by utilizing the reflection positivity and the chessboard estimate, to obtain an estimation of the partition function $Z^{\Lambda}$, which is given by,
$Z^{\Lambda}\geq(C_{1}\epsilon exp(2C_{2}\beta))^{|\Lambda|}\quad,$ (3)
where $\beta$ is the temperature, and they assume that $C_{1}=C_{2}\approx\cos(\pi/20)\approx 1$. To demonstrate the possibility of a first-order transition, they devised another partition function. This function considers a configuration divided into two regions: an ”ordered” region where the central site and its neighbors all reside within the ditch and a ”disordered” region where the central spin and its neighbors are not in the same ditch. They integrated over all configurations compatible with this contour, resulting in a partition function with a universal contour,$Z^{\Lambda}_{uc}$ given by:
$Z^{\Lambda}_{uc}\leq((2\epsilon)^{3/4}\exp(\beta))^{|\Lambda|}\quad.$ (4)
From the estimates of the partition functions, it is possible to demonstrate, using theorem 1 in the paper by Van Enter et al. [10], the coexistence of two Gibbs measures. This is achieved by considering
$\frac{Z^{\Lambda}_{uc}}{Z^{\Lambda}}\leq\epsilon^{\frac{|\Lambda|}{(4+C_{3})}}$ (5)
In essence, if $\epsilon$ is less than 1, $Z^{\Lambda}$ dominates the fraction, resulting in a suppression of the contours between the ordered and disordered sites, leading to a completely disordered system. For $\epsilon$ that exceeds 1, the partition function $Z^{\Lambda}_{uc}$ dominates this fraction and we have a ”ordered” system. Thus, it is possible to exist a temperature region where the two partition functions are comparable, implying the non-uniqueness of the Gibbs measures.
In the work of Van Enter et al. [15], they did not estimate the critical temperature or the critical parameter q where the first-order transition appears. Here, we define a function for $\epsilon$ that preserves the system’s interactions on average as a function of the parameter $q$. This function is given by a combination of gamma functions as follows:
$\epsilon(q)=\frac{\sqrt{\pi}}{2}\left(\frac{2^{-q}\sqrt{\pi}\Gamma[\frac{1}{2}+q]}{\Gamma\left[\frac{1+q}{2}\right]^{2}}\right)^{-1}\frac{\Gamma\left[\frac{1+q}{2}\right]}{\Gamma\left[1+\frac{q}{2}\right]}$ (6)
Applying the conditions for a first-order phase transition ($\epsilon\leq 1$), we obtained the critical parameter $q_{c}\geq 2.28851$. Since $q$ is an integer, the smallest critical parameter is equal to $q_{c}=3$, which coincides with the value obtained through simulations. The previous equation also allows us to estimate the transition temperature for any $q$ value. With this equation in hand, we can again divide eq.(4) by eq.(3) and, through a simple algebraic manipulation, as follows, obtain:
Table 1: Critical and transition temperature for different values of q by the methods discussed.
q FSS $\chi$ FSS $\Upsilon$ BKT $U_{L}$ First Order $U_{L}$ Gibbs Measures
1 $0.699\pm 0.001$ $0.692\pm 0.005$ $0.710\pm 0.001$------ - - - - - ------- - - - - - ------ - - - -
2 $0.679\pm 0.003$ $0.662\pm 0.005$ $0.688\pm 0.003$------ - - - - - ------- - - - - - ------ - - - -
3 $0.655\pm 0.001$ $0.638\pm 0.005$ $0.657\pm 0.005$ $0.682\pm 0.005$ $0.685\pm 0.001$ $0.68\pm 0.05$
4 $0.634\pm 0.001$ $0.621\pm 0.005$ $0.633\pm 0.003$ $0.662\pm 0.003$ $0.657\pm 0.001$ $0.65\pm 0.05$
5 $0.616\pm 0.001$ $0.608\pm 0.005$ $0.613\pm 0.004$ $0.640\pm 0.004$ $0.637\pm 0.001$ $0.64\pm 0.05$
6 ------- - - - - - $0.599\pm 0.005$------ - - - - - $0.623\pm 0.002$ $0.621\pm 0.001$ $0.62\pm 0.04$
7 ------- - - - - - $0.586\pm 0.005$------ - - - - - $0.608\pm 0.001$ $0.607\pm 0.001$ $0.61\pm 0.04$
$\frac{Z^{\Lambda}_{uc}}{Z^{\Lambda}}\leq\frac{|(2\epsilon)^{\frac{3}{4}}exp(\beta)|^{|\Lambda|}}{|C_{1}\epsilon exp(2C_{2}\beta)|^{|\Lambda|}}\leq 1$ (7)
$\exp[\beta(1-2C_{2})-\frac{1}{4}ln(\epsilon)+\frac{3}{4}ln(2)-ln(C_{1})]\leq 1$ (8)
$\beta(1-2C_{2})-\frac{1}{4}ln(\epsilon(q))+\frac{3}{4}ln(2)-ln(C_{1})\leq 0$ (9)
Thus, we can estimate the critical temperature at which the first-order phase transition occurs by letting $C_{1}$ and $C_{2}$ be free parameters, expecting them to be lower than $\cos(45^{\circ})$ in accordance with the analytical results of Van Enter et al. The obtained results are shown in Table 1 and are consistent with the transition temperature obtained by histogram and Binder Fourth Cumulant. The angles obtained for both $C_{1}$ and $C_{2}$ were close to $\sim 40^{\circ}$ regardless of the parameter $q$; a value higher than that was obtained in the simulations considering the entire lattice, which was approximately $\sim 52^{\circ}$ in the transition temperature. Notably, the value of $q$ doesn’t seem to affect the azimutal angle value.
IV conclusion
In this study, we employed Monte Carlo simulations and analytical analysis to investigate the 2D generalized XY model. Our findings reveal not only characteristics of a first-order phase transition for $q\geq 3$, but also the presence of three distinct regions characterized by the existence of phase transitions. In the first region ($q\leq 3$), a BKT regime is observed where the energy and the vortex density vary continuously, consistent with the usual XY model. In the second region ($63\leq q<$ 6), a coexistence of first-order transition and BKT transition occurs, but at different temperatures, with the first-order transition temperature always higher than the BKT transition. In the third region ($q\geq 6$), the first-order transition persists, but the analyses for the critical temperature of the BKT transition do not provide conclusive results. This suggests a possible change in the universality of the system and the potential extinction of the BKT phase transition.
Another important result is the impact of the vortex density on the phase transitions of the system. To illustrate, consider the histogram for $q=6$ at the transition temperature $T_{c}=0.6212$ and $L=96$. The two different phases of the system are represented by figures 4c and 4a. Figure 4a represents an ”Ordered Phase” with a lower vortex density, while figure 4c is a ”Disordered Phase” with a higher vortex density. The proliferation of vortices is identified as the cause of the first-order phase transition. To connect this result with the findings of Van Enter et al., consider the ”disordered” state, where two spins are in different ditches. The weak interaction between the spins increases the energy of the system, leading to the proliferation of vortices as shown in 4c. Conversely, when the spins belong to the same ditch, the energy decreases rapidly, resulting in a decrease in the number of vortices. Therefore, the first-order transition from an ”ordered” state to a ”disordered state” is caused by a proliferation of vortices in the system. It is essential to note that the quotation marks around ”ordered” and ”disordered” emphasize the absence of real order in the system. One phase is merely more ordered than the other, and there is no actual order in the system. Additionally, the magnetization remains null in agreement with the Theorem of Mermin-Wagner and Shlosmann, making it a first-order transition without symmetry breaking.
In summary, our study offers a comprehensive understanding of the properties and phase transition behavior of the generalized XY model. The results obtained clearly demonstrate the occurrence of first-order phase transitions and elucidate their underlying mechanism. The identified classes of phase transitions align with the analytical results, and the mechanism of this new phase transition is attributed to the proliferation of vortices in the system. The next step is to explore whether other models exhibiting first-order phase transitions, such as the Domany model, can be explained in a similar manner.
Acknowledgements.
The authors thank financial support from CAPES (Finance Code 001), CNPq, FAPEMIG, INCT/CNPq - Spintrônica e Nanoestruturas Magnéticas Avançadas (INCT-SpinNanoMag), and Rede Mineira de Nanomagnetismo/FAPEMIG.- Brazil. R.J.C.L. acknowledges A. Tirelli for the valuable discussions and M. Capone for all the support in developing and finalizing this work. R.J.C.L. acknowledges the financial support of the National Recovery and Resilience Plan PNRR MUR, PE0000023-NQSTI, financed by the European Union - Next Generation EU and the MUR Italian National Centre for High-Performance Computing, Big Data and Quantum Computing, grant number CN00000013, as well as the financial support from Simon’s Foundation.
Appendix A Binder’s Fourth Cumulant
An initial estimation of the critical temperature can be derived from the Binder’s Fourth Cumulant defined as:
$U_{L}=1-\frac{\left\langle\left(M_{x}^{2}+M_{y}^{2}\right)^{2}\right\rangle}{2\left\langle M_{x}^{2}+M_{y}^{2}\right\rangle}$ (10)
where $M_{x}$ and $M_{y}$ represent the in-plane magnetization components. For any system size, this quantity reaches $0.5$ in the low-temperature phase $(T\ll T_{c})$ and tends towards zero in the high-temperature phase $(T\gg T_{c})$. When measured at $T_{c}$, its value is approximately independent of the system size. Therefore, $T_{c}$ can be estimated by plotting $U_{L}$ vs $T$ for different system sizes and identifying the point of intersection. In a first-order transition, the Cumulant assumes negative values [23, 22].
As a example, consider the parameter $q=3$ as in Figure 5a. We can observe the two transitions, a region where the cumulant of binder assumes negative values with a minimum at $T=0.68(2)$ for $L=128$ represents a first order transition, and a crossing point in $T=0.65(7)$ supposedly represents a BKT phase transition.
For $q=1$ to $q=5$ we can observe the BKT transition in the point where the curves intersect, however, for $q=6$ and higher the curves overlaps in more then one point, so its not possible anymore to estimate the critical temperature this way. Its a reliable technique, but different methods for determining the $T_{c}$ was considered.
Appendix B finite size scaling of magnetic susceptibility
(a) $q=3$
(b) $q=6$
(c) q=6 Binder Cumulant
Figure 6: Effects of finite-size scaling of helicity modulus for two different values of $q$ is observed in (a) and (b). With the increase of $q$, the helicity jump becomes inclined and starts at lower temperatures.(c) show FFS of the helicity modulus, for q=1 to 5, the point where the two lines cross is the critical temperature expected. 1.
A more accurate estimate for the critical temperature for a BKT transition involves the finite-size scaling analysis of the in-plane susceptibility $\chi$. To obtain the susceptibility of a component $\alpha$, the magnetization fluctuations are calculated as:
$\chi^{\alpha\alpha}=\frac{1}{Nk_{b}T}\left(\langle M_{\alpha}^{2}\rangle-\langle M_{\alpha}\rangle^{2}\right)$ (11)
where $M_{\alpha}$ represents the value of the magnetization obtained by summing over the component $\alpha$ for all spins of the lattice. The in-plane susceptibility is the average of the planar component susceptibilities:
$\chi=\frac{1}{2}\left(\chi^{xx}+\chi^{yy}\right)$ (12)
The finite-size scaling of the in-plane susceptibility is derived assuming that a power-law scaling of the susceptibility [32] holds near and below $T_{c}$ for any value of q,
$\chi\propto L^{2-\eta}$ (13)
the exponent $\eta$ describes the long-distance behavior of in-plane correlations below $T_{c}$. For the XY model, the critical temperature is reached when $\eta=1/4$. Initially, we assume that $\eta$ is valid for any value of $q$, and the assumption is tested by the quality of the produced scaling. Thus, $\chi/L^{7/4}$ is plotted against $t$ for various system sizes. The common crossing point of the curves yields the critical temperature. For $q=1$, the expected result for the critical temperature of $T_{c}=0.69(9)$ is obtained. With an increase in the parameter $q$, the critical temperature decreases. For example, in Figure 5e, the critical temperature obtained was $t=0.65(5)$, and for $q=6$, where determining the critical temperature by this method becomes challenging, a region of crossing points is observed, as depicted in Figure 5f. This behavior persists for higher values of $q$. The interpretation drawn is that the assumption of the value of the scale parameter is no longer valid, indicating a change in the order of the phase transition.
Appendix C Helicity Modulus and Finite-Size Scalling
The Helicity is a measure of the change in dimensionless free energy due to an infinitesimal spin twist across the system along one coordinate,
$\Upsilon(T)=\frac{1}{N}\frac{\partial^{2}f}{\partial\Delta^{2}}$ (14)
taking an infinitesimal spin twist at the x component for the generalized XY-model:
$\begin{split}\Upsilon(T)=&-\frac{1}{2}\left\langle H\right\rangle\\ &-\frac{J}{Nk_{b}T}\left\langle\left[\sum_{\left\langle i,j\right\rangle}(\sin\theta_{i}\sin\theta_{j})^{q}\sin(\phi_{i}-\phi_{j})\hat{e}_{ij}\cdot\hat{x}\right]^{2}\right\rangle\end{split}$ (15)
According to the renormalization group, the helicity modulus in an infinite system jumps from zero to a finite value $\frac{2}{\pi}k_{B}T_{c}$ at the critical temperature. Therefore, we can obtain $T_{c}$ by plotting the helicity as a function of the temperature and locating the intersection with the straight line. Due to the dependence on the size of the lattice, better results can be obtained using finite-size scaling of the system. A useful scaling expression [37, 38, 39, 40] for BKT is given by:
$\frac{\pi\Upsilon}{2k_{B}T}=A(T)\left[1+\frac{1}{2\ln\left(\frac{L}{L_{o}}\right)}\right]$ (16)
where $A(T)$ and $L_{o}$ are fitting constants. This expression is exact at $T=T_{c}$, where $A(T_{c})=1$. By plotting $A(T)$ for $T$, we can find the temperature where $A(T)=1$. The results are expressed in Figure 6c. The critical temperature obtained for $q=1$ is $T_{c}=0.69(2)$, which is slightly smaller than the literature value of $T_{c}=0.699$ and obtained by other methods. With the increase of the parameter $q$, as in the other methods, the critical temperature also decreases. The difference here is that we can’t observe any change of behavior that indicates an extinction in the BKT transition or the emergence of a new class of phase transition.